A person notices a mild electric shock if the current along a path through the thumb and the index finger exceeds 80.0 μΑ. Calculate the maximum allowable voltage without shock across the thumb and the index finger with a dry-skin resistance of 4.21×105 Ω Submit Answer Tries 0/99 Calculate the maximum allowable voltage with a wet-skin resistance of 1700 Ω Tries 0/99 Submit Answer
Answer #1

here,

the maximum current , I = 80 uA = 80 * 10^-6 A

the resistance of dry-skin , Rd = 4.21 * 10^5 ohm

the maximum allowable voltage for dry-skin , Vd = I * Rd

Vd = 80 * 10^-6 * 4.21 * 10^5 V

Vd = 33.68 V

the resistance of wet-skin , Rw = 1700 ohm

the maximum allowable voltage for wet-skin , Vw = I * Rw

Vw = 80 * 10^-6 * 1700 V

Vw = 0.136 V